# Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions

### v1

@@ III Discussion
Identifying Individual Residences and Workplaces in Shijiazhuang. A user’s home location is defined as the place where they spend the longest time between 8 PM and 7 AM, and the total duration of stay each night must exceed 6 hours. If these criteria are not met, the location of the home for that individual on that day is considered unknown. The workplace is defined as the location outside the individual’s home, where a he/she spends the most time between 7 AM and 8 PM, and the total duration of stay each day must exceed 2 hours. If these requirements are not met, the workplace data for that individual on that day is empty. Given five weekdays of data, a maximum of five home location assessments can be made for one individual. To mitigate the impact of base station spacing, base stations with home locations less than 250 meters apart are merged, using the location where the longest stay occurred as a substitute. Then, the individual’s nighttime data is updated to obtain the position of his/her home each night. For an individual with five days of home location assessments, the location is considered a clear residential characteristic only if at least four days point to the same place. Similarly, there are five records for the workplace. We define the workplace for an individual as the location that appears most frequently among the five assessments. If two locations occur with the same number of times, the total duration of stays over the five days at both locations is compared, with the longer duration determining the individual’s workplace.
@@
The inferred home-departure time is the average time each commuter first leaves home between 5 AM and 11 AM on the 5 weekdays covered by the data. Commuters not meeting this criterion are assigned departure times proportionally. As shown in Fig. S2 of the SI, the proportion of commuters departing home between 7 AM and 8 AM is the highest. The home-departure time after home swapping is inferred as the time the commuter has to start their journey in order to reach their workplace at the same time before home swapping. The distributions of home-departure time after GHS and SDGHS are shown in Fig. S3A and S3B of the SI.
Average Congestion Coefficient. To explore the impact of shortened commuting distances on traffic congestion after home-swapping, we define an average congestion coefficient $I_{t}$ as follows:
@@
where $N_{\alpha}^{t}$ is the number of vehicles traveling in the same direction on road segment $\alpha$ at time $t$, and we calculate the start time of commuters on each segment based on their travel times and departure times. $\tilde{N}_{\alpha}$ is the maximum traffic volume of the road $\alpha$, given by
@@
which varies depending on the speed limit $\tilde{v}_{\alpha}$ of road segment $\alpha$(see navigation data in SI Section 1). $l_{c}$ is the average length of a vehicle, i.e. 4 meters; $l_{a}^{\alpha}$ is the length of the safety distance between vehicles, which is dependent on the maximum driving speed of road $\alpha$, e.g. $l_{a}^{\alpha}=100m$ if speed limit $\tilde{v}_{\alpha}=100km/h$ mcshane1990traffic .
The higher the average congestion coefficient, the more congested the city is. It is also important to note that the coefficient is calculated hourly from data in four 15-minute intervals, and commuters using non-motorized vehicles have been excluded. If a commuter passes road $\alpha$ during time interval $t$, then $N_{\alpha}$ at time $t$ increased by one.
Identifying Housing Price. We identify the price of a housing unit at a residence location based on their price listed in the second-hand market from the website Lianjia.com for Shijiazhuang, where information such as the name of the properties, the latitude and longitude of the residential communities, their total price and the unit price, are available (see SI Section S1). The city was divided into a grid of $500\times 500$ squares, and residences in the second-hand market, i.e. houses, apartments, etc., were placed in the corresponding grid based on their latitude and longitude. The average housing price of a residence in each grid is calculated by averaging the prices of all the second-hand residences within the grid. For grids without housing price data, the price of a housing unit is assumed to be the average of the housing prices of the nearest four grids.
@@
$D_{xy}=1-\frac{\vec{A}_{x}\cdot\vec{A}_{y}}{|\vec{A}_{x}||\vec{A}_{y}|}$ (9)
Urban Carbon Dioxide Emissions For exploring how GHS affects carbon dioxide emissions from the whole city, we define the carbon dioxide emission $E(t)$ during each time interval $t$ as in kissinger2019detailed :
@@
where $N_{c}$ is the number of commuters traveling at time interval $t$; $Path_{i}(t)$ is the set of road segments commuter $i$ passes during time interval $t$, and $d_{i}^{\alpha}$ is the length of road segment $\alpha$; $\zeta$ is the carbon dioxide emission factor, which is equal to 0.209 $kgCO_{2}e/km$; $c_{\alpha}$ is the number of vehicles passing road $\alpha$ at the same time with commuter $i$; $S_{\alpha}$ and $R_{\alpha}$ represent the slope and the type of road $\alpha$, which are both set as constants for simplicity. In addition to the total carbon emissions, we can also obtain the emissions from each commuter or on each road in each time interval.
@@ Acknowledgments.
C.Z., C.H.Y. and A.Z. designed the research, C.Z., Y.L., X.H. and J.D. performed the experiments, C. Z., C.H.Y. and A.Z. analyzed the data, C.H.Y. and A.Z. wrote the paper.
@@
Data and materials availability. The raw data cannot be shared by the authors due to the Mobile Privacy Policy of China, but might be available upon reasonable request submitted to the mobile network operator (China Mobile) at http://it.10086.cn/services/##\##dsjArea.
@@
Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions
Chen Zhao, Yuqing Liu, Xiaoyue Hou, Jianghui Ding, Chi Ho Yeung, and An Zeng
@@ IV Data Description
  • Shenzhen Living-Working Dataset [2]: The Shenzhen data is a dataset collected by Zhilu Yuan and others, representing the largest urban living-working tracking dataset from the Shenzhen 2015 Census logs. It includes records of living and working locations for 5 million anonymous commuters, covering 71 geographically adjacent subdivisions across 10 districts. The locations of each commuters’ residence and workplace are indicated by the latitude and longitude of their respective streets.
  • Abidjan Call-Detail Record (CDR) Dataset [3]: The Cote d’Ivoire CDR data provides anonymous user IDs, timestamps, and sub-county IDs (including their locations). When a user makes a call from a new location, the user’s movement can be identified.
@@
  • Shijiazhuang Commuting Route Navigation Data: To obtain the actual commuting distances before and after commuters swap their homes, first, different residential and workplace locations are paired. Due to the large volume of data, in this study we aggregate different location points to the nearest road center. If the residence and workplace are on the same road, the original location points are used. In this study, we used the aggregated locations to crawl the driving routes for each pair of commuting locations on the Baidu map [4], during morning traffic peak (6:00 A.M.-10:00 A.M.) and early evening traffic peak (4:00 P.M.-8:00 P.M.) in weekdays, which includes the time taken for an individual to travel through each road segment, the type and the distance of the road segment, and navigation details. The final data include the individual’s starting location, the aggregated starting location, commuting distance, commuting time, and commuting route.
  • Shenzhen Commuting Route Navigation Data: The residential-workplace dataset of Shenzhen is a street-level record, and we represent the whole street using the road center’s latitude and longitude, and crawl the driving routes for each pair of the commuting locations on the Baidu map.
@@
  • Shijiazhuang On-demand Vehicle Dataset: The data from DiDi on-demand vehicle in Shijiazhuang for five weekdays in July 2017, including the pick-up and drop-off location of each trip.
@@
  • Shijiazhuang Housing Price Data: The information of second-hand housing prices from Lianjia.com [5] within the Shijiazhuang city area in 2017, including total price, unit price, community name, and community location.
  • Shenzhen Housing Price Data: The information of second-hand housing prices from Lianjia.com [6] and Anjuke.com [7] within the Shenzhen city area in 2015, including total price, unit price, community name, and community location.
@@
  • Shijiazhuang Point-of-interest (POI) Data: The information on various types of Points of Interest (POI) within the Shijiazhuang city area collected from Gaode in 2017, including store name, type, location, area, and street information.
  • Shenzhen POI Data: Information on various types of Points of Interest (POI) within the Shenzhen city area collected from Gaode in 2015, including store name, type, location, area, and street information.
  • Shijiazhuang 2017 Census Data and Shenzhen 2015 Census Data [8].
@@
  • Shijiazhuang Urban Road Network Data: Obtained from OpenStreetMap [9] to facilitate aggregating residence or workplace locations to the nearest road center. The total number of road segment is 28.9 million. The average street length in the city, calculated from crawling the length of streets at all levels, is found to be 272 meters.
@@ V Statistical Analysis Based on Data
Identifying Commuters from the Same Household. To obtain the actual number of commuter households from the data, we take reference of the 7-th Census of Shijiazhuang, revealing the proportion of households from one-person to six-person to be 0.12, 0.27, 0.23, 0.19, 0.097, and 0.093 respectively. According to the demographic yearbook, one commuter can support up to 1.92 people, hence 422,454 commuters can support a maximum of $N_{p}=811112$ people. Denoting $p_{i}$ is the proportion of households with $i$ members, the number of households with $i$ members is $n_{i}=p_{i}\times N$, where $N$ is the total number of households. The number one-person to six-person households are found to be 30,870, 69,458, 59,168, 48,878, 24,953, and 23,924 respectively.
Next, we assumed that the number of commuters in one household is just the supply minimum $n_{c}^{min}$, or $n_{c}^{min}-1$, without considering supply redundancy. Therefore, the potential number of commuters in households with different number of members is shown in the following table:
@@
Table 1: The possible commuter families in different household types.
@@
Commuter Households Obtained through Trajectories Similarity. Since individual mobility trajectory data do not include the information about household membership, we used data from the 4 days in the 2 weekends to identify commuters from the same household. According to the census information, most of households in Shijiazhuang have less than $7$ members, and each working person supports up to 1.92 individuals. Therefore, we assume that each household has a maximum of 6 members and 4 commuters. We identify family members by calculating the similarity between the mobility trajectories among individuals who share the same residence location. By analyzing the distribution of similarities $S_{ij}$ between individuals, those with $S_{ij}>0.5$ are considered household members. An individual might have similarities over 0.5 with multiple people; these similarities are sorted from the highest to the lowest, and only a maximum of five individuals with the highest similarity are considered to be potential household members. Ultimately, the ratio of single-, double-, triple- and quadruple-commuter households was found to be 65:26:6:2.
@@
Compared to our census-deduced household ratios of 37:31:14:1, double and triple-commuter household are under-estimated, while single and quadruple-commuter households are over-estimated. To match the constructed and the deduced household ratio, we randomly selected 3,302 quadruple-commuter households, using the three members with the highest similarity in each household to form triple-commuter household, with the remaining person form a single-commuter household, resulting in 192,777 single-, 76,533 double-, 21,529 triple-, and 3,006 quadruple-commuter households. In this case, the number of single-commuter households still exceeded the inferred amounts, due to the criteria of $S_{ij}\geq 0.5$, which is a more stringent criteria for identifying house members, and thus leads to an overestimation of single-commuter households.
@@
To identify households with a ratio in agreement with the census of Shijiazhuang, we release the stringent criteria and form households among individuals from single-commuter households with $S_{ij}\geq 0.1$. We finally obtained a ratio of households matching the deduced ratio, such that single to quadruple-commuter household to 108,515, 90,843, 40,076, and 3,006 respectively.
@@
Using the shortest travel distance by motor vehicle. The mode of transport is highly related to the commuting distance, we analyzed the distance for on-demand motor vehicles the Didi in China in 2017 (as shown in Figure S1)to identify a threshold for urban residents to differentiate between short and long distances. As we can see in Fig. S1, the number of on-demand travel peaks at a distance of 2.5 km. Therefore, we assume that 2.5 km is the shortest distance that urban residents opt for motor vehicle travel in their daily commuting travel between their home and workplace.
@@
Empirical Commuter Departure Times.
@@
Commuters’ empirical departure time is set as the average time of their department from their home between 5 AM and 11 AM on the 5 weekdays in the dataset. The return time from the workplace for commuters is set as the average time of their departure from their workplace between 4 PM and 11 PM on the 5 weekdays in the dataset, as shown in Figure S2. Commuters departures are notably observed during the morning rush hour between 7-8 AM (Fig. S2A), and most commuters returns are observed during the evening rush hour between 5-6 PM (Fig. S2B), though the return time distribution is noticeably more dispersed.

### v2

@@ III Discussion
Identifying Individual Residences and Workplaces in Shijiazhuang. A user’s home location is defined as the place where they spend the longest time between 8 p.m. and 7 a.m., and the total duration of stay each night must exceed 6 hours. If these criteria are not met, the location of the home for that individual on that day is considered unknown. The workplace is defined as the location outside the individual’s home, where he/she spends the most time between 7 a.m. and 8 p.m., and the total duration of stay each day must exceed 2 hours. If these requirements are not met, the workplace data for that individual on that day is empty. Given five weekdays of data, a maximum of five home location assessments can be made for one individual. To mitigate the impact of base station spacing, base stations with home locations less than 250 meters apart are merged, using the location where the longest stay occurred as a substitute. Then, the individual’s nighttime data is updated to obtain the position of his/her home each night. For an individual with five days of home location assessments, the location is considered a clear residential characteristic only if at least four days point to the same place. Similarly, there are five records for the workplace. We define the workplace for an individual as the location that appears most frequently among the five assessments. If two locations occur with the same number of times, the total duration of stays over the five days at both locations is compared, with the longer duration determining the individual’s workplace.
@@
The inferred home-departure time is the average time each commuter first leaves home between 5 a.m. and 11 a.m. on the 5 weekdays covered by the data. Commuters not meeting this criterion are assigned departure time proportionally. As shown in Fig. S2 of the SI, the proportion of commuters departing home between 7 a.m. and 8 a.m. is the highest. The home-departure time after home swapping is inferred as the time the commuter has to start their journey in order to reach their workplace at the same time before home swapping. The distributions of home-departure time after GHS and SDGHS are shown in Fig. S3A and S3B of the SI.
Average Congestion Coefficient. To explore the impact of shortened commuting distances on traffic congestion after home-swapping, we define an average congestion coefficient $I_{t}$ as follows
@@
where $N_{\alpha}^{t}$ is the number of vehicles traveling in the same direction on road segment $\alpha$ at time interval $t$, such that we calculate the time of commuters starting on each segment based on their travel time and departure time. It is important to note that the time point is calculated hourly collected from four 15-minute intervals, and commuters using non-motorized vehicles have been excluded. If a commuter passed road $\alpha$ during the time interval $t$, the volume of the road $\alpha$ at time interval $t$ is incremented by one. Hence, the real motor vehicle volume on every road in each time interval is accurately counted. It is stipulated that the morning arrival time at workplace and evening departure time from workplace remain unchanged. Thus, after GHS, the morning departure time and evening arrival time from the new residence can be deduced from navigation data, such that the new motor vehicle volume on every road can be estimated in any time interval. $\tilde{N}_{\alpha}$ is the maximum traffic volume of the road $\alpha$, given by
@@
which varies depending on the speed limit $\tilde{v}_{\alpha}$ of road segment $\alpha$(see navigation data in SI Section 1). $l_{c}$ is the average length of a vehicle, i.e. 4 meters; $l_{a}^{\alpha}$ is the length of the safety distance between vehicles, which is dependent on the maximum driving speed of road $\alpha$, e.g. $l_{a}^{\alpha}=100m$ if speed limit $\tilde{v}_{\alpha}=100km/h$ mcshane1990traffic . In consideration of the impact of multi-modal transportation, the city’s entire bus network was integrated. The average length of automobiles is $l_{c}$, whilst the average length of buses is four times of that of automobiles.
Identifying Housing Price. We identify the price of a housing unit at a residence location based on their price listed in the second-hand market from the website Lianjia.com for Shijiazhuang, where information such as the name of the properties, the latitude and longitude of the residential communities, their total price and the unit price, are available (see SI Section S1). We have discretized the city into $0.005^{\circ}$ ($\sim 0.5km\times 0.5km$) grid cells and identified the housing price /sq.m/sq.m averaged over all transacted residences for each individual cell, capturing localized housing costs while accounting for spatial heterogeneity. For grids without housing price data, the price of a housing unit is assumed to be the average of the housing prices of the nearest four grids.
@@
$D_{xy}=1-\frac{\vec{A}_{x}\cdot\vec{A}_{y}}{|\vec{A}_{x}||\vec{A}_{y}|},$ (11)
Urban Carbon Dioxide Emissions For exploring how GHS affects carbon dioxide emissions from the whole city, we define the carbon dioxide emission $E(t)$ during each time interval $t$ as in ref. kissinger2019detailed
@@
where $N_{c}$ is the number of commuters traveling at time interval $t$; $Path_{i}(t)$ is the set of road segments commuter $i$ passes during time interval $t$, and $d_{i}^{\alpha}$ is the length of road segment $\alpha$; $\zeta$ is the carbon dioxide emission factor zhou2022identifying , which is equal to 0.209 $kgCO_{2}e/km$; $c_{\alpha}$ is the number of vehicles passing road $\alpha$ at the same time with commuter $i$; $S_{\alpha}$ and $R_{\alpha}$ represent the slope and the type of road $\alpha$, which are both set as constants for simplicity. In addition to the total carbon emissions, we can also obtain the emissions from each commuter or on each road in each time interval. Additionally, given that the buses use diesel, the carbon emission factor is 0.281 $kgCO_{2}e/km$.
Data-driven model of residential choice. We construct the data-driven model of residential choice based on households’ preference as following
$A_{i}^{j}\propto\frac{E_{i}^{j}\times U_{i}}{D_{i}^{j}},$ (13)
here, $E_{i}^{j}=1-\frac{\left|B^{j}-R_{i}\right|}{B^{j}}$. $R_{i}$ represents the rental price of location $i$, which is proportional to the housing price of location $i$; $B^{j}$ represents the observed rental expenditure of household $j$ (economic capacity proxy). Residential choice probability follows an inverse exponential relationship with price disparity $\left|B^{j}-R_{i}\right|$, reflecting households’ preference for residence matching their financial capacity.
In Henderson’s urban system model, there exists an inverted-U shaped relationship between regional scale and residents’ regional utility henderson2014economic . Based on this theory, we assume that the attractiveness of a residential location is proportional to the utility value of that region. The utility of location $i$ for household $j$ is related to the population size of the region: when the population size of location $i$ is small, all resident households can benefit; but as the population size increases, the per capita accessibility to infrastructure amenities begins to decline, and consequently, the probability of household $j$ in choosing location $i$ as residence decreases. Therefore, we define the impact of location $i$’s locational utility ($U_{i}$) on household $j$’s choice as $\sqrt{(S_{i}-O_{i})(O_{i}+k)}$, such that $S_{i}$ represents the total population capacity of location $i$ as a residential area, $O_{i}$ represents the current population size of location $i$ and $k$ is the initial parameter, which is set to $2$ in our model.
The attractiveness of area $i$ for household $j$ is formulated as follows
$A_{i}^{j}\propto\left(1-\frac{\left|B^{j}-R_{i}\right|}{B^{j}}\right)\left(\frac{\sqrt{(S_{i}-O_{i})(k+O_{i})}}{\sqrt[m]{{\textstyle\prod_{p=1}^{m}d_{iW}^{C^{j}_{p}}}}}\right),$ (14)
where $d_{iW}^{C^{j}_{p}}$ represents the commuting distance of the $p$-th commuting members in household $j$ if they reside at location $i$, and $m$ denotes the number of commuting individuals in the household. To balance the influence of each family member’s commuting distance, we characterize this using the geometric mean of family members’ commuting distances, i.e., $D_{i}^{j}=\sqrt[m]{{\textstyle\prod_{p=1}^{m}d_{iW}^{C^{j}_{p}}}}$. To prevent division by zero, the minimum commuting distance for any family member is set to $1m$ in the simulation.
Considering the restriction in living space, all the households could be divided into two groups: single/double-commuter households, and triple/quadruple-commuter households. Households in the same group can exchange their residences, but not between groups, i.e., 1-commuter households and 2-commuter households can exchange their residences, as well as between 3-commuter and 4-commuter households. As the number of households of each group at location $i$ remains unchanged, the total number of residents may be more than the empirical capacity of location $i$. Therefore, $S_{i}$ can be defined as the numerical upper bound, which is equal to the sum of twice the number of single/double-commuter households and four times the number of triple/quadruple-commuter households.
The G-residence is defined as the residential location determined for each household through the GHS algorithm, exhibiting significant heterogeneity across different households. Consequently, the subsidy also varies substantially among households, as formalized in the following theoretical expression
$\alpha_{i}^{j}\propto\left[1-\frac{\theta}{(1+\frac{d_{iG}}{L})^{2}}\right]^{-\delta},$ (15)
For household $j$, $\alpha_{i}^{j}$ represents the incentive provision by government policy, where location $g$ is the government-recommended G-residence for household $j$ (as shown in Fig. 8A), and $\theta$ denotes the economic subsidy rate (ranging from $0\%$ to $100\%$) that household $j$ would receive when residing at location $g$. To allow feasibility in this incentive scheme, we incorporate a distance-decay mechanism for the subsidy, where $L$ is the average street distance in the target city and $d_{iG}$ is the distance between location $i$ and the G-residence. The smaller $d_{iG}$, the closer the subsidy amount at location $i$ to $\theta$ for household $j$. The parameter $\delta$ represents the household’s sensitivity to subsidy incentives - higher values indicate more responsive behavior from household $j$ to the policy incentives. In our simulations, this parameter is set to $\delta=4$.
@@ Acknowledgments.
Y.-Q.L., C.Z., C.H.Y. and A.Z. designed the research, Y.-Q.L., C.Z. , X.-Y.Y., X.H. performed the experiments, Y.-Q.L., C.Z., X.-Y.Y., C.H.Y. and A.Z. analyzed the data, C.Z., C.H.Y. and A.Z. wrote the paper.
@@
Data and materials availability. The raw data cannot be shared by the authors due to the Mobile Privacy Policy of China, but might be available upon reasonable request submitted to the mobile network operator (China Mobile) at http://it.10086.cn/services/#\#dsjArea.
@@
Guiding Self-Organizing Dynamics of Residential Choice in Cities to Reduce Traffic Congestion and Carbon Emissions
Yu-Qing Liu, Chen Zhao, Xiao-Yong Yan, Xiaoyue Hou, Chi Ho Yeung, An Zeng
@@ IV Data Description
  • Shijiazhuang Commuting Route Navigation Data: To obtain the actual commuting distances before and after commuters swap their homes, first, different residential and workplace locations are paired. Due to the large volume of data, in this study we aggregate different location points to the nearest road center. If the residence and workplace are on the same road, the original location points are used. In this study, we used the aggregated locations to crawl the driving routes for each pair of commuting locations on the Baidu map [2], during the morning traffic peak (6:00 A.M.-10:00 A.M.) and early evening traffic peak (4:00 P.M.-8:00 P.M.) in weekdays, which includes the time taken for an individual to travel through each road segment, the type and the distance of the road segment, and navigation details. The final data include the individual’s starting location, the aggregated starting location, commuting distance, commuting time, and commuting route.
@@
  • Shijiazhuang On-demand Vehicle Dataset: The data from DiDi on-demand vehicles in Shijiazhuang for five weekdays in July 2017, including the pick-up and drop-off location of each trip.
@@
  • Shijiazhuang Housing Price Data: The information on second-hand housing prices from Lianjia.com [3] within the Shijiazhuang city area in 2017, including total price, unit price, community name, and community location.
@@
  • Shijiazhuang Point-of-interest (POI) Data [8]: The information on various types of Points of Interest (POI) within the Shijiazhuang city area collected from Gaode Map in 2017, including store name, type, location, area, and street information.
@@
  • Shijiazhuang Urban Road Network Data: Obtained from OpenStreetMap [4] to facilitate aggregating residence or workplace locations to the nearest road center. The total number of road segments is 28.9 million. The average street length in the city, calculated from crawling the length of streets at all levels, is found to be 272 meters.
  • Shijiazhuang bus route data: All bus routes in Shijiazhuang were crawled from the public transportation website [5], and the total number of bus routes is 246, and the total number of bus stops is 3661. The data include bus name, bus ID, starting station, terminal station, and passing bus station.
  • Shijiazhuang bus departure schedule data [5]: The data include bus name, starting station, terminal station, first and last departure time.
  • Urban Mobility Dataset of Chinese Major Cities [6]: The aggregated and anonymised mobile phone dataset was provided by a Chinese telecommunications operator. The raw dataset consists of mobile phone signal records collected over a two-month period, representing approximately 90 million residents across 60 cities. The number of anonymised users represents, on average, $18\%$ of the total population and corresponds to the proportion of mobile phone users to the overall population across the 60 cities studied. The operator partitioned all cities into a 1 km $\times$ 1 km grid cell. The aggregated, non-personally identifiable human movement records represent the number of trips between grids per hour in August and November 2019 in 60 Chinese cities. For characterizing the commuting behavior of residents, using data only covering the time span between 7:00 and 7:59, which corresponds to the typical morning peak hour.
  • Chinese Major Cities’ Commuting Route Navigation Data: Residence and workplace pairs were first defined by linking trip origins and destinations. Due to the 1 km grid resolution, the precise latitude and longitude for each location were represented by the coordinates of the grid center. Driving routes for each pair were then collected from the Baidu Map API [2]. The collected data include travel time for each road segment, road type, segment distance, and detailed navigation information. The final dataset is structured identically to that of Shijiazhuang.
  • Chinese Major Cities’ Housing Price Data [7]: The information of second-hand housing prices from national residential estate historical transactions in 2019, including total price, unit price, community name, and community location.
  • Chinese Major Cities’ POI Data [8]: The information on various types of POIs within China’s 28 major cities was collected from Gaode Map in 2019, including store name, type, location, area, and street information.
@@ V Statistical Analysis Based on Data
Identifying commuters from the same household. To obtain the actual number of commuter households from the data, we take reference of the 7-th Census of Shijiazhuang [9], revealing the proportion of households from one-person to six-person to be 0.12, 0.27, 0.23, 0.19, 0.097, and 0.093 respectively. According to the demographic yearbook [10], one commuter can support up to 1.92 people, hence 422,454 commuters can support a maximum of $N_{p}=811112$ people. Denoting $p_{i}$ is the proportion of households with $i$ members, the number of households with $i$ members is $n_{i}=p_{i}\times N$, where $N$ is the total number of households. The number one-person to six-person households are found to be 30,870, 69,458, 59,168, 48,878, 24,953, and 23,924 respectively.
Next, we assumed that the number of commuters in one household is just the supply minimum $n_{c}^{min}$, or $n_{c}^{min}-1$, without considering supply redundancy. Therefore, the potential number of commuters in households with different number of members is shown in the Supplementary Table S1.
@@
Supplementary Table S1. The possible commuter families in different household types.
@@
Commuter Households Obtained through Trajectories Similarity. Since individual mobility trajectory data do not include the information about household membership, we used data from the 4 days in the 2 weekends to identify commuters from the same household. According to the census information, most of households in Shijiazhuang have less than $7$ members, and each working person supports up to 1.92 individuals. Therefore, we assume that each household has a maximum of 6 members and 4 commuters. We identify family members by calculating the similarity between the mobility trajectories among individuals who share the same residence location. By analyzing the distribution of similarities $S_{ij}$ between individuals, those with $S_{ij}\geq 0.5$ are considered household members. An individual might have similarities over 0.5 with multiple people; these similarities are sorted from the highest to the lowest, and only a maximum of five individuals with the highest similarity are considered to be potential household members. Ultimately, the ratio of single-, double-, triple- and quadruple-commuter households was found to be 65:26:6:2.
@@
Compared to our census-deduced household ratios of 37:31:14:1, double and triple-commuter households are under-estimated, while single and quadruple-commuter households are over-estimated. To match the constructed and the deduced household ratio, we randomly selected 3,302 quadruple-commuter households. In each selected quadruple-commuter household, we use the three members with the highest similarity in each household to form a triple-commuter household, with the remaining person forming the single-commuter household, resulting in 192,777 single-, 76,533 double-, 21,529 triple-, and 3,006 quadruple-commuter households. In this case, the number of single-commuter households still exceeded the inferred amounts, due to the criteria of $S_{ij}\geq 0.5$, which is a more stringent criteria for identifying household members, and thus leads to an overestimation of single-commuter households.
@@
To identify households with a ratio in agreement with the census of Shijiazhuang, we release the stringent criteria and form households among individuals from single-commuter households with $S_{ij}\geq 0.1$. We finally obtained a ratio of households matching the deduced ratio, such that single to quadruple-commuter households to 108,515, 90,843, 40,076, and 3,006 respectively.
@@
Using the shortest travel distance by motor vehicle. The mode of transport is highly related to the commuting distance, we analyzed the distance for on-demand motor vehicles the Didi in China in 2017 (as shown in Figure S1) to identify a threshold for urban residents to differentiate between short and long distances. As we can see in Figure S1, the number of on-demand travel peaks at a distance of 2.5 km. Therefore, we assume that 2.5 km is the shortest distance that urban residents opt for motor vehicle travel in their daily commuting travel between their home and workplace. This finding is in alignment with the 2020 National Commuting Report of China, in the classification of non-motorized vs. motorized travel [11].
@@
Empirical commuter departure time.
@@
Commuters’ empirical departure time is set as the average time of their department from their home between 5 AM and 11 AM on the 5 weekdays in the dataset. The return time from the workplace for commuters is set as the average time of their departure from their workplace between 4 PM and 11 PM on the 5 weekdays in the dataset, as shown in Figure S2. Commuters departures are notably observed during the morning rush hour between 7-8 AM (Figure S2A), and most commuters returns are observed during the evening rush hour between 5-6 PM (Figure S2B), though the return time distribution is noticeably more dispersed.